OpenStudy (anonymous):

Gretchen says she is thinking of a number. She says if she adds 9 to her number and multiplies by 3, the result is 72. What is Gretchen’s number? a. 15 b. 24 c. 5 d. 33

OpenStudy (anonymous):

please help

OpenStudy (dumbcow):

3(n+9) = 72 n+9 = 24 n = 15

OpenStudy (anonymous):

Answer is A. You could use each answer to figure it out. 15 works because 15+9=24 And 24x3=72 |

OpenStudy (anonymous):

At the end of a board game, Al had 57 game dollars. During the game he had won $200, lost $150, won $25, lost $10, and lost $35. How much money did Al have at the start? a. $47 c. $177 b. $27 d. $107

OpenStudy (anonymous):

helppppppppppppppppppppppppp